Give an example of how a fast-food restaurant, or a car wash, or any other company of your choice can use economies of scope to achieve cost savings. Then give an example of how the same company can use economies of scale to save money.
Answer #1

Answer:

Example of a fast food restaurant economies of scope-

The economic concept that refers to the decrease in the total cost of production when a range of products are produced together rather than separately.

Formula for Economies of Scope- (C(qa) + C(qb) - C (qa+qb)) / C (qa+qb)

  • C(qa) =cost of producing quantity qa of good a separately
  • C(qb) =cost of producing quantity qb of good b separately
  • C(qa+qb) = cost of producing quantities qa and qb
  • S = percentage cost saving when the goods are produced together.
  • Therefore, S would be greater than 0 when economies of scope exist.

example:

A restaurant produces both hamburgers and sandwiches. The cost of separately producing 1,000,000 hamburgers is 0.50 each. Likewise, if 4,000,000 sandwiches are produced separately, the cost is 0.30 each. If 1,000,000 hamburgers and 4,000,000 sandwiches are produced together (by using the same preparation and storage facility), the total cost is 1,500,000.

To determine the economies of scope:

  1. Determine C(qa) = 1,000,000 * 0.50 = 500,000
  2. Determine C(qb) = 4,000,000 * 0.30 = 1,200,000
  3. Determine C(qa+qb) = 1,500,000
  4. Plug the numbers into the Economies of Scope formula

(500,000 + 1,200,000 – $1,500,000) / 1,500,000 = 13.33%. Therefore, the cost of producing hamburgers and sandwiches together is 13.33% less than the cost of producing them separately.

Economies of scale to save money -

taking bulk orders can save money for a restaurant. Fast-food chains take advantage of bulk food orders as ordering in bulk makes the raw material more cheap.